Crew Photos – April 13th

Photo of the Day: Yuri’s Night Fun

We donned our crew shirts, Yuri’s Night pins, and used some space-y holographic face masks to celebrate Yuri Gagarin, the first man in space.

Operations Report – April 13th

Crew 192 Operations Report 12Apr2018


Name of person filing report: Richard Blakeman

Non-nominal systems: 1. Refrigerator (resolved); 2. Kitchen sink faucet

Notes on non-nominal systems: The refrigerator appeared to have stopped working overnight. Outlet was inspected, fridge unplugged and plugged back in, has been working since. Will continue to monitor.

Sink is being worked on right now.

Generator (hours run): 13 hours

Solar ‒ SOC 98%

Diesel – 75% fuel

Propane – 66% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – Shannon told us we were not to worry about this gas. We have no knowledge of the gas or that 5 gallon container.

Water (trailer) – Empty, tank being cleaned in RAM.
Water (static) – 925 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133428.9

Toilet tank emptied: No. Last emptied April 11, 2018.

ATV’s Used: Blue ATVs #1 and #3

Oil Added? No but checked and OK.

ATV Fuel Used: 10%

# Hours the ATVs were Used today: 3

Notes on ATVs: All working well and ready for use.

Deimos rover used: No
Hours: 119.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: No
Hours: 31
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Opportunity rover used: No
Hours: 29.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: No
Hours: 25.2
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

HabCar used and why, where? No

General notes and comments: Crew is doing well.

Summary of internet: 0MB at 1900.

Summary of suits and radios: Suits all charging, 2 one piece suits used. Radios working fine.

Summary of Hab operations: Cleaned and organized hab, including ceiling lights in the upper deck and various bins and shelves in lower deck. Organized food cabinets and cleaned additional kitchen appliances and pans. Fixed kitchen sink faucet.

Summary of GreenHab operations: See GreenHab report.

Summary of ScienceDome operations: None

Summary of RAM operations: The trailer water tank remains in the RAM for additional cleaning.

Summary of health and safety issues: None

Questions, concerns, and requests to Mission Support:

– Regarding use of tables in the lower deck: There is a 3 foot wooden table in the spacesuit room and two 5 foot plastic tables in the lower deck. Requesting permission to swap the wooden table with one of the longer plastic ones so we can accommodate all one-piece EVA suits in the spacesuit room.

– Request permission to change out water filter that was delivered earlier today.

– Reviewed possible fixes for the skirt of the HAB as requested. Suggest using sheet metal screws and the electric drill to fasten skirting in place. Respectfully request permission to perform this repair via engineering EVA.

– There’s a picture attached of the base of a small tv screen/computer monitor that is in the hab, but does not seem to have the proper cords available for use. Does anyone know if the cords are available anywhere in the hab site?

– The laptop labeled “Hab” in the upper deck also does not have the proper cords to check for power. Please advise.

Thank you for removing the unneeded items we left outside. There are a few additional items we put outside for pickup.

EVA Report – April 13th

EVA Report

EVA #7 – 12Apr2018 – Sol 5 ‒ Cancelled due to mix-up in schedule

Crew members: Joe Dituri, Ashok Narayanamoorthi, Victoria Varone





EVA Report

EVA #8 – 13Apr2018 – Sol 5

Crew members: Richard Blakeman, Andreea Radulescu

– Check out Kissing Camels site
– RAM activities upon return
– Hab survey upon return

1500 – 1800

Explored the region and confirmed ability to establish base comms when at the top of the ridge. Found evidence of missing crew member and produced for discussion with team. Proceeded into the RAM upon return to further clean the water tank and then looked at the skirt of the HAB to find out what is required to repair.

– Blue ATV #1
– Blue ATV #3

Sol Summary – April 13th

Crew 192 Sol Summary Report 13Apr2018

Sol 5

Summary Title: The Winds of Mars

Author: Victoria Varone

Mission Status: Our Mars mission has been slightly inhibited by approximately two and a half days of strong winds (paging Mark Watney), but like most scientists, the crew is keeping themselves busy and happy finding other things to do (repairs, cleaning, organization, cooking) until the weather calms down.

Sol Activity Summary: With strong winds blowing again today, the crew attempted to gather a weather report before heading out on their first EVA of the day. Unfortunately, a sluggish connection prevented us from obtaining this prior to the EVA (Martian technology, right?). We left a little later than planned but were instructed to call off the outing, so we headed back inside to make a new plan for the morning and afternoon.

The winds started to pick up again a bit, so the crew used the extra time inside to conduct a day of cleaning and organizing the hab. This turned out to be a worthwhile task as we removed layers of dust from many areas and long-unused objects and organized some of the boxes and bins in the lower deck. After this, the crew enjoyed a delicious brunch of nachos crafted by their own Richard Blakeman.

We went ahead with our afternoon EVA as scheduled as the winds were slower than they were yesterday and our two crew members remained perfectly safe during their outing. Upon their return, we had a nice dinner, checked out the GreenHab together, looked through some of our many amazing mission photos so far, and planned for weekend EVAs and activities.

Look Ahead Plan: The forecast says that the winds will die down tonight and the weather will be beautiful over the weekend. We’ll use that time to make up for the EVAs and other activities that were prohibited over the last couple of days due to the strong winds.

Anomalies in Work: None

Weather: Chilly and windy. Winds blowing at approximately 23 mph with gusts 30 mph.

Crew Physical Status: All crew are physically healthy and safe.

EVA: See EVA Report.

Reports to be filed: Sol Summary Report, GreenHab Report, EVA Request, EVA Report Journalist Report, Photo Report

Support Requested: See Operations Report

Maintenance request: See Operations Report

Greenhab Report – April 13th

GreenHab Report
Name: Andreea Radulescu
Date: 13 Apr 2018

Environmental control:

Ambient with door open
Shade cloth: on
Fan: not on.

Average temperatures: 23℃
Low 14.6℃
High 53.2℃

Hours of supplemental light: None

Changes to crops: None

Daily water usage for crops: 8 gal

Time(s) of watering for crops: Twice: In the morning around 10:15 am and afternoon 4:30pm

Moringa research observations: None

Changes to research plants: No

Aquaponics: None

Narrative: None

Support/supplies needed: None

Commander Report – April 13th

Happy Friday the 13th CAPCOM….Today we had two scheduled EVAs. We aborted one EVA due to a miscommunication on our part with mission control. Communication delays and network issues will always hamper us on Mars …we will persevere amd continue efforts to improve coms always.

The crew remains in good spirits and in my opinion a BUSY crew is a HAPPY crew. Since our EVA was aborted we took the opportunity to field day and straighten up the HAB. We organized the tool box placing tools correctly as written on outside of draws. (I know…a bit anal on our part…but everything has a place….) A few items of gear adrift were removed from the HAB and placed in pickup spot for mission control as directed by mission control.

We recieved answers to our request for maintenance and we will T/S the sink hot water as well as attempt a repair to the skirt on HAB. We removed and cleaned the moths / bugs from all upper deck light fixtures. It is so bright now…;*) We want to do all we can to improve the state of this living facility.

IMPRESSED: I am so impressed with our new call sign assignment of “Poison Ivy” for Andreea (Greenhab officer – Blue 03). She brought the garden back from a dry less productuve state (due to water issues last rotation) to a garden that continues to produce many small tomatoes despite our ability to consume them daily. BZ Poison Ivy….

LEARNED I learned that the internet is not reliable and I need to be much earlier to be proactive about weather.

POA&M We intend to perform two sets of EVAs tomorrow and if approved continue training crew and repairs to HAB. We may even be able to finish cleaning the water tank in the RAM and return to service for next crew. If approved we will also perform skirt repairs at end of EVA 2.

~Dituri sends

Warm Regards,

Joseph Dituri, PhD(c)
CDR, U.S. Navy Diving Officer (ret)
Director – IBUM

Crew Photos – April 12th

Journalist Report – 12 April 2018

Author: Victoria Varone

Today was one of those days that highlights just how passionate, adaptable, and hard-working space dreamers really are.

In the face of weather that disrupted our EVA plans, the whole crew stepped up to the plate, offering up an unbelievable variety of skills and ideas to make repairs and adjustments with a limited supply of tools and hardware. They pooled their backgrounds, areas of expertise, physical capabilities, and improvisation skills to come up with a solution to the problem at hand.

Nothing brings out the strength of such dreamers more than a problem that needs to be solved, be it an Earthly issue or a Mars mission. I’m incredibly proud of this crew today and never except anything less than amazing when we’re together.

In honor of Yuri’s Night: “When I orbited the Earth in a spaceship, I saw for the first time how beautiful our planet is. Mankind, let us preserve and increase this beauty, and not destroy it!”

GreenHab Report – 12 April 2018

Name: Andreea Radulescu
Date: 12 Apr 2018

Environmental control:

Ambient with door open
Shade cloth on
Fan not on.

Average temperatures: 39.7℃
Low 14.2℃
High 53.2℃

Hours of supplemental light: None

Changes to crops: Was able to pick a lot of tomatoes.

Daily water usage for crops: 5 gal

Time(s) of watering for crops: Once in the morning around 10:00 am

Moringa research observations: None

Changes to research plants: No

Aquaponics: None

Narrative: Picked few more tomatoes for the team which we will consume at dinner.

Support/supplies needed: None

EVA Report – 12 April 2018

EVA #6 – 12Apr2018 – Sol 4

Crew members: Andreaa Radulescu, Ashok Narayanamoorthi, Richard Blakeman

Scout for good sites to perform medical emergency simulations.

0845 Start EVA
0900 Surveyed URC North Site
0930 Returned to HAB at request of commander
1000 EVA end, transfer to engineering EVA.

This EVA was meant for scouting sites for a medical emergency situation. One site was surveyed before the crew members on the EVA were recalled to help with a repair in the HAB from damage caused by severe winds.

1. Blue #1
2. Blue #2
3. Blue #3

EVA #7 – 12Apr2018 – Sol 4

Crew members: Richard Blakeman, Ashok Narayanamoorthi, Joe Dituri, Andreaa Radulescu, Victoria Varone

– All crew transfer to the RAM and clean and sanitize the water tank.
– Garbage pails tied to trailer as they were blowing away in the strong wind.

1630 – 1745

Today’s second EVA was a productive engineering EVA. The tank is 85% cleaned and we intend to complete further final cleaning tomorrow.

– None